# Dealing with unbounded gradients in stochastic saddle-point optimization

### v1

@@ 3.2 Sub-bilinear games and general divergences
Suppose that $f$ is sub-bilinear and the stochastic gradients satisfy the conditions in Equation (4). Letting $\varrho_{x}=\frac{\eta_{y}L^{2}}{\gamma_{y}}$ and $\varrho_{y}=\frac{\eta_{x}L^{2}}{\gamma_{x}}$, and $\bm{x}^{*},\bm{y}^{*}$ be arbitrary comparator points, the expected duality gap of COMIDA satisfies the following bound:
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\frac{\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)}{\eta_{y}T}+\frac{\varrho_{y}{\left\|\bm{y}^{*}-\bm{y}_{1}\right\|^{2}_{x,*}}}{2}$
$\quad+\frac{\mathcal{D}_{x}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)}{\eta_{x}T}+\frac{\varrho_{x}{\left\|\bm{x}^{*}-\bm{x}_{1}\right\|^{2}_{y,*}}}{2}$
$\quad+L^{2}\left(\frac{\eta_{y}}{2\gamma_{y}}+\frac{\eta_{x}}{2\gamma_{x}}\right).$
The most important special case of our setting is when the norms appearing in the statement are dual to each other, and in particular $*\left\|\cdot\right\|_{x}=\left\|\cdot\right\|_{y,*}$ and $*\left\|\cdot\right\|_{y}=\left\|\cdot\right\|_{x,*}$, so that $\omega_{x}$ is strongly convex with respect to $*\left\|\cdot\right\|_{y,*}$ and $\omega_{y}$ is strongly convex with respect to $*\left\|\cdot\right\|_{x,*}$. This is the case for instance when $\mathcal{X}=\mathcal{Y}=^{m}$, $\omega_{x}=\frac{1}{2}\left\|\cdot\right\|_{\bm{A}}^{2}$ and $\omega_{y}=\frac{1}{2}\left\|\cdot\right\|_{\bm{A}^{-1}}^{2}$ for a symmetric positive definite matrix $\bm{A}\in^{m\times m}$. We state a specialized version of our statement to this setting below.
@@
Suppose that $f$ is sub-bilinear and the stochastic gradients satisfy the conditions in Equation (4), and suppose additionally that $\omega_{x}$ is $\gamma_{x}$-strongly convex with respect to $*\left\|\cdot\right\|_{y,*}$ and $\omega_{y}$ is $\gamma_{y}$-strongly convex with respect to $*\left\|\cdot\right\|_{x,*}$. Set the parameters as $\varrho_{x}=\frac{\eta_{y}L^{2}}{\gamma_{y}}$, $\varrho_{y}=\frac{\eta_{x}L^{2}}{\gamma_{x}}$, $\eta_{x}=\eta_{y}=\sqrt{\gamma_{x}\gamma_{y}/L^{2}T}$. Then, letting $\bm{x}^{*},\bm{y}^{*}$ be arbitrary comparator points, the duality gap of COMIDA satisfies the following bound:
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]=\mathcal{O}\left(\frac{L\left(\mathcal{D}_{x}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)+\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)+1\right)}{\sqrt{\gamma_{x}\gamma_{y}T}}\right).$
@@ 4 Application to Average Reward Markov Decision Processes
$\mathbb{E}\left[\left\langle\bm{\mu}^{\pi^{*}}-\bm{\mu}^{\overline{\pi}_{T}},\bm{r}\right\rangle\right]\leq\frac{\mathcal{D}_{\textup{KL}}\left(\bm{\mu}^{\pi^{*}}\middle\|\bm{\mu}_{1}\right)}{\eta_{\mu}T}+\eta_{\mu}+4\eta_{v}$
@@
Thus, the iteration complexity of COMIDA-MDP for finding an $\varepsilon$-optimal policy is of the order $\frac{B^{4}SA\log(SA)}{\varepsilon^{2}}$. As each iteration uses $SA+1$ queries to the generative model, this makes for a total of $\frac{B^{4}S^{2}A^{2}\log(SA)}{\varepsilon^{2}}$ query complexity, which is suboptimal in terms of its dependence on $SA$, but optimal in terms of $\varepsilon$. Most importantly, this guarantee constitutes the first one we are aware of in the literature that does not require prior knowledge of the so-called “bias span” $B$.
@@ 5 Discussion
We leave several interesting questions open for future work. The biggest of these questions is if the scaling with the initialization error $\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}+\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}$ can be improved to $\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}+\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}$. This is obviously possible when we have prior knowledge of these norms, and can tune the learning rate to fully optimize the first set of bounds in Theorem 3.1. Without prior knowledge, it is less clear if such improvement is possible, unlike in the case of convex minimization problems where there exist efficient algorithms that achieve such improved rates, at least up to log factors (Streeter & McMahan, 2012; Orabona, 2013, 2014). Another interesting line of investigation is to find out if it is possible to extend our methodology to go significantly beyond bilinear objectives.
We close by recalling that our approach bears some significant similarity with the stabilized online mirror descent method of Fang et al. (2022): their approach introduces a similar regularization term to address issues faced by OMD in unconstrained convex minimization problems. This idea was adapted to equilibrium finding in multiplayer games by Hsieh et al. (2021), but their results are once again not comparable to ours even in two-player zero-sum games (e.g., they consider noiseless gradients and bounded decision sets). We are curious to see if this stabilization trick can find further uses in the context of saddle-point optimization and game theory in the future.
@@ Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 950180).
@@ Appendix A Proof of results in Section 3
To control the first set of terms in the above expression, we apply the regret analysis of COMIDA in Section A.2.1. Precisely, with $\mathcal{D}_{x}\!\left(\bm{x}\middle\|\bm{x}^{\prime}\right)=\frac{1}{2}\left\|\bm{x}-\bm{x}^{\prime}\right\|^{2}_{2}$, $\mathcal{D}_{y}\!\left(\bm{y}\middle\|\bm{y}^{\prime}\right)=\frac{1}{2}\left\|\bm{y}-\bm{y}^{\prime}\right\|^{2}_{2}$ and, $\mathcal{H}_{x}(\bm{x})=\frac{1}{2}\left\|\bm{x}-\bm{x}_{1}\right\|^{2}_{2}$, $\mathcal{H}_{y}(\bm{y})=\frac{1}{2}\left\|\bm{y}-\bm{y}_{1}\right\|^{2}_{2}$, we get:
@@
$\qquad\qquad\qquad\leq\frac{\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}}{2\eta_{y}}+\frac{\eta_{y}}{2}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{y}(t)\right\|_{2}^{2}\right]+\frac{\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}}{2\eta_{x}T}+\frac{\eta_{x}}{2}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{x}(t)\right\|_{2}^{2}\right].$ (6)
@@
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\left(\frac{1}{2\eta_{y}T}+\frac{\varrho_{y}}{2}\right)\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}+\frac{\eta_{y}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)^{\mathsf{\scriptscriptstyle T}}\bm{x}_{1}-\widehat{\bm{c}}(t)\right\|_{2}^{2}\right]$
$\quad+\left(\frac{1}{2\eta_{x}T}+\frac{\varrho_{x}}{2}\right)\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}+\frac{\eta_{x}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)\bm{y}_{1}+\widehat{\bm{b}}(t)\right\|_{2}^{2}\right]$
$\quad+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\bm{y}_{t}-\bm{y}_{1}\right\|_{2}^{2}\right]\left(\eta_{x}L_{M}^{2}-\frac{\varrho_{y}}{2}\right)+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\bm{x}_{t}-\bm{x}_{1}\right\|_{2}^{2}\right]\left(\eta_{y}L_{M}^{2}-\frac{\varrho_{x}}{2}\right).$
By setting $\varrho_{y}=2\eta_{x}L_{M}^{2}$ and $\varrho_{x}=2\eta_{y}L_{M}^{2}$, we eliminate the last two terms in the bound above and arrive at the result stated in the theorem:
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\left(\frac{1}{2\eta_{y}T}+\eta_{x}L_{M}^{2}\right)\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}+\frac{\eta_{y}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)^{\mathsf{\scriptscriptstyle T}}\bm{x}_{1}-\widehat{\bm{c}}(t)\right\|_{2}^{2}\right]$
$\quad+\left(\frac{1}{2\eta_{x}T}+\eta_{y}L_{M}^{2}\right)\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}+\frac{\eta_{x}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)\bm{y}_{1}+\widehat{\bm{b}}(t)\right\|_{2}^{2}\right].$
@@
Consider the expected duality gap at arbitrary comparator points $(\bm{x}^{*},\bm{y}^{*})$:
@@ A.2.1 Regret Analysis of COMIDA on a Regularized Objective
This part of the proof is based on the regret analysis of Composite Mirror Descent (COMID) for stochastic convex optimization. The proof is a more-or-less standard exercise in convex analysis (appearing, e.g., as Theorem 8 of Duchi et al. (2010)), and we provide it for completeness as Lemma D.1 in Appendix D. In this section, we directly apply the implied guarantee on the regret of COMID against a fixed comparator in Corollary D.2 to control the regret of each player.
@@
Notice that $\ell_{t}^{\left(\text{reg}\right)}\left(\cdot\right)$ is convex by the concave property of $f(\bm{x}_{t},\cdot)$. We will bound the regret using Corollary D.2, with initial iterate $\bm{u}_{1}=\bm{y}_{1}$, gradient estimates $\widetilde{\bm{{g}}}_{u}(t)=-\widetilde{\bm{{g}}}_{y}(t)$ and gradients $\bm{g}_{\bm{u}}(t)=-\bm{g}_{y}(t)$. Also, we will set $\mathcal{U}=\mathbb{R}^{n}$, $\omega_{u}=\omega_{y}$, $\eta_{u}=\eta_{y}$ and $\varrho_{u}=\varrho_{y}$. With $\bm{y}^{*}$ fixed and independent of the iterates, this gives
$\sum_{t=1}^{T}\mathbb{E}\left[f^{\left(\text{reg}\right)}\left(\bm{x}_{t},\bm{y}^{*}\right)-f^{\left(\text{reg}\right)}\left(\bm{x}_{t},\bm{y}_{t}\right)\right]\leq\frac{\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)}{\eta_{y}}+\frac{\eta_{y}}{2\gamma_{y}}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{y}(t)\right\|_{y,*}^{2}\right]+\varrho_{y}\mathbb{E}\left[\mathcal{H}_{y}(\bm{y}_{1})\right].$
Likewise, reusing previous notation we denote the loss of the min player as in round $t$ as $\ell_{t}(\bm{x})=f(\bm{x},\bm{y}_{t})$. Since $f(\cdot,\bm{y}_{t})$ is convex, and by equivalence of the minimization step of COMIDA to that of Corollary D.2 when $\bm{u}_{1}=\bm{x}_{1}$, $\widetilde{\bm{{g}}}_{u}(t)=\widetilde{\bm{{g}}}_{x}(t)$, $\bm{g}_{\bm{u}}(t)=\bm{g}_{x}(t)$, $\mathcal{U}=\mathbb{R}^{m}$, $\omega_{u}=\omega_{x}$, $\eta_{u}=\eta_{x}$ and $\varrho_{u}=\varrho_{x}$, we can bound the regret of the min player against a fixed comparator $\bm{x}^{*}$ as follows:
$\sum_{t=1}^{T}\mathbb{E}\left[f^{\left(\text{reg}\right)}\left(\bm{x}_{t},\bm{y}_{t}\right)-f^{\left(\text{reg}\right)}\left(\bm{x}^{*},\bm{y}_{t}\right)\right]\leq\frac{\mathcal{D}_{x}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)}{\eta_{x}}+\frac{\eta_{x}}{2\gamma_{x}}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{x}(t)\right\|_{x,*}^{2}\right]+\varrho_{x}\mathbb{E}\left[\mathcal{H}_{x}(\bm{x}_{1})\right].$
@@
$\quad\leq\frac{\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)}{\eta_{y}}+\frac{\eta_{y}}{2\gamma_{y}}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{y}(t)\right\|_{y,*}^{2}\right]+\varrho_{y}\mathbb{E}\left[\mathcal{H}_{y}(\bm{y}_{1})\right]$
$\quad\qquad+\frac{\mathcal{D}_{x}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)}{\eta_{x}}+\frac{\eta_{x}}{2\gamma_{x}}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{x}(t)\right\|_{x,*}^{2}\right]+\varrho_{x}\mathbb{E}\left[\mathcal{H}_{x}(\bm{x}_{1})\right].$

### v2

@@ 3.2 Sub-bilinear games and general divergences
Suppose that $f$ is sub-bilinear and the stochastic gradients satisfy the conditions in Equation (4). Letting $\varrho_{x}=\frac{2\eta_{y}L^{2}}{\gamma_{y}}$ and $\varrho_{y}=\frac{2\eta_{x}L^{2}}{\gamma_{x}}$, and $\bm{x}^{*},\bm{y}^{*}$ be arbitrary adaptive comparator points, the expected duality gap of COMIDA satisfies the following bound:
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\frac{2\mathbb{E}\left[\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)\right]}{\eta_{y}T}+\frac{\varrho_{y}}{2}\mathbb{E}\left[\left\|\bm{y}^{*}-\bm{y}_{1}\right\|^{2}_{x,*}\right]$
$\quad+\frac{2\mathbb{E}\left[\mathcal{D}_{x}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)\right]}{\eta_{x}T}+\frac{\varrho_{x}}{2}\mathbb{E}\left[\left\|\bm{x}^{*}-\bm{x}_{1}\right\|^{2}_{y,*}\right]$
$\quad+L^{2}\left(\frac{\eta_{y}}{\gamma_{y}}+\frac{\eta_{x}}{\gamma_{x}}\right).$
The proof is provided in Appendix A.2. Notably, as in the case of Theorem 3.1, the statement remains valid for adaptively chosen comparator points such as $(\bm{x}^{*},\bm{y}^{*})=\mathop{\mbox{ arg\,max}}_{\bm{x},\bm{y}\in\mathcal{S}}G(\bm{x},\bm{y})$ for arbitrary bounded sets $\mathcal{S}\subseteq\mathcal{X}\times\mathcal{Y}$. The most important special case of our setting is when the norms appearing in the statement are dual to each other, and in particular $\left\|\cdot\right\|_{x}=\left\|\cdot\right\|_{y,*}$ and $\left\|\cdot\right\|_{y}=\left\|\cdot\right\|_{x,*}$, so that $\omega_{x}$ is strongly convex with respect to $\left\|\cdot\right\|_{y,*}$ and $\omega_{y}$ is strongly convex with respect to $\left\|\cdot\right\|_{x,*}$. This is the case for instance when $\mathcal{X}=\mathcal{Y}=^{m}$, $\omega_{x}=\frac{1}{2}\left\|\cdot\right\|_{\bm{A}}^{2}$ and $\omega_{y}=\frac{1}{2}\left\|\cdot\right\|_{\bm{A}^{-1}}^{2}$ for a symmetric positive definite matrix $\bm{A}\in^{m\times m}$. We state a specialized version of our statement to this setting below.
@@
Suppose that $f$ is sub-bilinear and the stochastic gradients satisfy the conditions in Equation (4), and suppose additionally that $\omega_{x}$ is $\gamma_{x}$-strongly convex with respect to $\left\|\cdot\right\|_{y,*}$ and $\omega_{y}$ is $\gamma_{y}$-strongly convex with respect to $\left\|\cdot\right\|_{x,*}$. Set the parameters as $\varrho_{x}=\frac{2\eta_{y}L^{2}}{\gamma_{y}}$, $\varrho_{y}=\frac{2\eta_{x}L^{2}}{\gamma_{x}}$, $\eta_{x}=\eta_{y}=\sqrt{\gamma_{x}\gamma_{y}/L^{2}T}$. Then, letting $\bm{x}^{*},\bm{y}^{*}$ be arbitrary adaptive comparator points, the duality gap of COMIDA satisfies the following bound:
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$
$\qquad=\mathcal{O}\left(\frac{L\left(\mathbb{E}\left[\mathcal{D}_{x}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)+\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)\right]+1\right)}{\sqrt{\gamma_{x}\gamma_{y}T}}\right).$
@@ 4 Application to Average Reward Markov Decision Processes
$\mathbb{E}\left[\left\langle\bm{\mu}^{\pi^{*}}-\bm{\mu}^{\overline{\pi}_{T}},\bm{r}\right\rangle\right]\leq\frac{\mathcal{D}_{\textup{KL}}\left(\bm{\mu}^{\pi^{*}}\middle\|\bm{\mu}_{1}\right)}{\eta_{\mu}T}+\eta_{\mu}+2\eta_{v}$
@@
Thus, the iteration complexity of COMIDA-MDP for finding an $\varepsilon$-optimal policy is of the order $\frac{B^{4}SA\log(SA)}{\varepsilon^{2}}$. We stress, unlike similar prior results such as the ones of Wang (2017); Jin & Sidford (2020); Cheng et al. (2020) that this result does not require prior knowledge of $B$. As each iteration uses $SA+1$ queries to the generative model, this makes for a total of $\frac{B^{4}S^{2}A^{2}\log(SA)}{\varepsilon^{2}}$ query complexity, which is suboptimal in terms of its dependence on $SA$, but optimal in terms of $\varepsilon$.
@@ 5 Discussion
We leave several interesting questions open for future work. The biggest of these questions is if the scaling with the initialization error $\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}+\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}$ can be improved to $\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}+\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}$. This is obviously possible when we have prior knowledge of these norms, and can tune the learning rate to fully optimize the first set of bounds in Theorem 3.1. Without prior knowledge, it is less clear if such improvement is possible, unlike in the case of convex minimization problems where there exist efficient algorithms that achieve such improved rates, at least up to log factors (Streeter & McMahan, 2012; Orabona, 2013, 2014). While we were not aware of this at the time of preparing the first version of this article, the recent work of Jacobsen & Cutkosky (2023) has already provided some results that hint at a negative answer: their Theorem 2.3 shows that there exists an online linear optimization problem with sub-quadratic gradient growth where quadratic scaling with the comparator norm is unavoidable. Whether or not their counterexample can be adapted to our setting remains to be seen.
We close by highlighting (one more time) some similarities between our approach and some previously proposed methods. We first mention the work of Jacobsen & Cutkosky (2023) that we learned about after completing the first version of the present manuscript. Their methods have the advantage of being completely parameter-free, and demonstrating a slightly more refined dependence on the comparator norms than our guarantees do. On the other hand, our algorithm is arguably much simpler and is thus much easier to adapt to more general settings, as evidenced by our main results that are stated in terms of general Bregman divergences. In contrast, the analysis of Jacobsen & Cutkosky (2023) is strictly tied to Euclidean norms and it is unclear if a generalization to other geometries is straightforwardly possible. Similarly, we believe that adjusting their analysis to handle stochastic gradients and data-dependent comparators may not be entirely straightforward. Besides this work, our method also bears close similarity to the stabilized online mirror descent method of Fang et al. (2022): their approach introduces a similar regularization term to address issues faced by OMD in unconstrained convex minimization problems. Their use of regularization had the purpose of allowing time-dependent (and more generally, adaptive) learning-rate schedules, which is ultimately quite different from the purpose that we employed this technique for, and also requires a different tuning rule than our method. The extension of this technique by Hsieh et al. (2021) to a two-player game setting similar to ours remained restricted to consider noiseless gradients and bounded decision sets. Additionally, an anonymous reviewer has pointed our attention to the similarity between our approach and an “anchoring” technique extensively studied under the name of Halpern iteration in the optimization literature (Halpern, 1967; Lieder, 2021). All these connections suggest that the simple and natural regularization trick we made use of in this paper is a tool of fundamental importance with a large range of diverse uses. In light of our results and these observations, we are particularly curious to see if this trick will find further uses in the context of saddle-point optimization and game theory in the future.
@@ Acknowledgements
The authors wish to thank the exceptionally dedicated anonymous reviewer that pointed out and helped us fix an issue with our definition of the duality gap in the original version of the paper. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 950180).
@@ Appendix A Proof of results in Section 3
To control the first set of terms in the above expression, we apply the regret analysis in Section A.2.1. Precisely, with $\mathcal{D}_{x}\!\left(\bm{x}\middle\|\bm{x}^{\prime}\right)=\frac{1}{2}\left\|\bm{x}-\bm{x}^{\prime}\right\|^{2}_{2}$, $\mathcal{D}_{y}\!\left(\bm{y}\middle\|\bm{y}^{\prime}\right)=\frac{1}{2}\left\|\bm{y}-\bm{y}^{\prime}\right\|^{2}_{2}$ and, $\mathcal{H}_{x}(\bm{x})=\frac{1}{2}\left\|\bm{x}-\bm{x}_{1}\right\|^{2}_{2}$, $\mathcal{H}_{y}(\bm{y})=\frac{1}{2}\left\|\bm{y}-\bm{y}_{1}\right\|^{2}_{2}$, we get:
@@
$\qquad\qquad\qquad\leq\frac{\mathbb{E}\left[\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}\right]}{\eta_{y}}+\eta_{y}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{y}(t)\right\|_{2}^{2}\right]+\frac{\mathbb{E}\left[\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}\right]}{\eta_{x}T}+\eta_{x}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{x}(t)\right\|_{2}^{2}\right].$ (6)
@@
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\left(\frac{1}{\eta_{y}T}+\frac{\varrho_{y}}{2}\right)\mathbb{E}\left[\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}\right]+\frac{2\eta_{y}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)^{\mathsf{\scriptscriptstyle T}}\bm{x}_{1}-\widehat{\bm{c}}(t)\right\|_{2}^{2}\right]$
$\quad+\left(\frac{1}{\eta_{x}T}+\frac{\varrho_{x}}{2}\right)\mathbb{E}\left[\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}\right]+\frac{2\eta_{x}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)\bm{y}_{1}+\widehat{\bm{b}}(t)\right\|_{2}^{2}\right]$
$\quad+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\bm{y}_{t}-\bm{y}_{1}\right\|_{2}^{2}\right]\left(2\eta_{x}L_{M}^{2}-\frac{\varrho_{y}}{2}\right)+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\bm{x}_{t}-\bm{x}_{1}\right\|_{2}^{2}\right]\left(2\eta_{y}L_{M}^{2}-\frac{\varrho_{x}}{2}\right).$
By setting $\varrho_{y}=4\eta_{x}L_{M}^{2}$ and $\varrho_{x}=4\eta_{y}L_{M}^{2}$, we eliminate the last two terms in the bound above and arrive at the result stated in the theorem:
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\left(\frac{1}{\eta_{y}T}+2\eta_{x}L_{M}^{2}\right)\mathbb{E}\left[\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}\right]+\frac{2\eta_{y}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)^{\mathsf{\scriptscriptstyle T}}\bm{x}_{1}-\widehat{\bm{c}}(t)\right\|_{2}^{2}\right]$
$\quad+\left(\frac{1}{\eta_{x}T}+2\eta_{y}L_{M}^{2}\right)\mathbb{E}\left[\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}\right]+\frac{2\eta_{x}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)\bm{y}_{1}+\widehat{\bm{b}}(t)\right\|_{2}^{2}\right].$
@@
Consider the expected duality gap at arbitrary adaptive comparator points $(\bm{x}^{*},\bm{y}^{*})$:
@@ A.2.1 Regret Analysis of COMIDA on a Regularized Objective
This part of the proof is based on the regret analysis of Composite Mirror Descent (COMID) for stochastic convex optimization. The proof is a more-or-less standard exercise in convex analysis (appearing, e.g., as Theorem 8 of Duchi et al. (2010)), and we provide it for completeness as Lemma D.1 in Appendix D. In this section, we directly apply the implied guarantee on the regret of COMID against an adaptive comparator in Corollary D.2 to control the regret of each player.
@@
Notice that $\ell_{t}^{\left(\text{reg}\right)}\left(\cdot\right)$ is convex by the concave property of $f(\bm{x}_{t},\cdot)$. We will bound the regret using Corollary D.2, with initial iterate $\bm{u}_{1}=\bm{y}_{1}$, gradient estimates $\widetilde{\bm{{g}}}_{u}(t)=-\widetilde{\bm{{g}}}_{y}(t)$ and gradients $\bm{g}_{\bm{u}}(t)=-\bm{g}_{y}(t)$. Also, we will set $\mathcal{U}=\mathbb{R}^{n}$, $\omega_{u}=\omega_{y}$, $\eta_{u}=\eta_{y}$ and $\varrho_{u}=\varrho_{y}$. With $\bm{y}^{*}$ adaptive and potentially dependent on the interaction history, this gives
$\sum_{t=1}^{T}\mathbb{E}\left[f^{\left(\text{reg}\right)}\left(\bm{x}_{t},\bm{y}^{*}\right)-f^{\left(\text{reg}\right)}\left(\bm{x}_{t},\bm{y}_{t}\right)\right]\leq\frac{2\mathbb{E}\left[\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)\right]}{\eta_{y}}+\frac{\eta_{y}}{\gamma_{y}}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{y}(t)\right\|_{y,*}^{2}\right]+\varrho_{y}\mathbb{E}\left[\mathcal{H}_{y}(\bm{y}_{1})\right].$
Likewise, reusing previous notation we denote the loss of the min player as in round $t$ as $\ell_{t}(\bm{x})=f(\bm{x},\bm{y}_{t})$. Since $f(\cdot,\bm{y}_{t})$ is convex, and by equivalence of the minimization step of COMIDA to that of Corollary D.2 when $\bm{u}_{1}=\bm{x}_{1}$, $\widetilde{\bm{{g}}}_{u}(t)=\widetilde{\bm{{g}}}_{x}(t)$, $\bm{g}_{\bm{u}}(t)=\bm{g}_{x}(t)$, $\mathcal{U}=\mathbb{R}^{m}$, $\omega_{u}=\omega_{x}$, $\eta_{u}=\eta_{x}$ and $\varrho_{u}=\varrho_{x}$, we can bound the regret of the min player against an adaptive comparator $\bm{x}^{*}$ as follows:
$\sum_{t=1}^{T}\mathbb{E}\left[f^{\left(\text{reg}\right)}\left(\bm{x}_{t},\bm{y}_{t}\right)-f^{\left(\text{reg}\right)}\left(\bm{x}^{*},\bm{y}_{t}\right)\right]\leq\frac{2\mathbb{E}\left[\mathcal{D}_{x}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)\right]}{\eta_{x}}+\frac{\eta_{x}}{\gamma_{x}}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{x}(t)\right\|_{x,*}^{2}\right]+\varrho_{x}\mathbb{E}\left[\mathcal{H}_{x}(\bm{x}_{1})\right].$
@@
$\quad\leq\frac{2\mathbb{E}\left[\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)\right]}{\eta_{y}}+\frac{\eta_{y}}{\gamma_{y}}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{y}(t)\right\|_{y,*}^{2}\right]+\varrho_{y}\mathbb{E}\left[\mathcal{H}_{y}(\bm{y}_{1})\right]$
$\quad\qquad+\frac{2\mathbb{E}\left[\mathcal{D}_{x}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)\right]}{\eta_{x}}+\frac{\eta_{x}}{\gamma_{x}}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{x}(t)\right\|_{x,*}^{2}\right]+\varrho_{x}\mathbb{E}\left[\mathcal{H}_{x}(\bm{x}_{1})\right].$